OpenStudy (ttp):

A movie rental store sells a couples movie pack for six dollars and a family movie pack for ten dollars. In one year, the Nguyen family paid one hundred and twenty-six dollars for 13 movie packs. How many family movie packs did they pay for? 9 10 11 12

OpenStudy (anonymous):

umm..... tricky

OpenStudy (ttp):

I know :)

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

oh well now I know why its tricky for me im in 8th grade lol

OpenStudy (ttp):

Lol

OpenStudy (anonymous):

but I will try im advanced

OpenStudy (ttp):

Ok, Thanks! :)

OpenStudy (lexipoo1998):

It is 9

OpenStudy (lexipoo1998):

oh wait no its not give me a second

OpenStudy (lexipoo1998):

it is actually 12

OpenStudy (lexipoo1998):

this is because when you multiply you get 12*10 = 12 then 1*6 = 6 120 + 6 = 126 and 12 + 1 = 13

OpenStudy (lexipoo1998):

Does that make sense?

OpenStudy (ttp):

Yes :)

OpenStudy (lexipoo1998):

I really hopoe you didn't just look at my first answer and put it because if you think about it if it was 9 which is was not you woud have 9*10 = 90 but then you would need 6 of the other packs to make 126 and that wuldnot add up to 13

OpenStudy (lexipoo1998):

Ok good :)

OpenStudy (ttp):

Thank you! :)

OpenStudy (lexipoo1998):

You're Welcome